When the current in the wire is 0.20 A, the magnetic field strength at the upper section of wire due to the lower section of wire is 1.3 x 10^-4 T.

Calculate the magnetic force acting per unit length on the upper section of wire.
1
Expert's answer
2019-10-22T10:26:09-0400
"\frac{F}{l}=BI=(1.3\cdot 10^{-4})(0.2)=2.6\cdot 10^{-5}\frac{N}{m}"